OpenStudy (anonymous):

hi i need help wtih this equation fidn the value of b satisfying the description the line through the points (1,5) and (-3,b) with the slope -2/7 the value of B is?

OpenStudy (chaise):

m= y-y/x-x -2/7 = b-5/-3-1 -2/7=b-5/-3-1 -2/7=b-5/-4 8/7=b-5 b=43/7 ~ 6.1 I hope this is right? I've never encountered a problem like this before.